# RBI-ThPhys-2024-15

### v1

@@ VI Modelling interference effects using reweighting
There are numerous examples of extended Higgs sector models that include an additional scalar that can impact the di-Higgs cross-section and differential distributions, see e.g. wg3ext for an overview of models currently investigated by the LHC experiments. The final di-Higgs spectrum depends on the parameters of the model. For example, in our investigated singlet model there are three free parameters that affect the relative contributions of the diagrams shown in Fig. 2 to the total cross-section, and there are countless other models that may contain even more free parameters. While in principle it is possible to generate Monte Carlo (MC) events for different combinations of parameters, and simulate the detector responses, the computational resources required would be immense. We have therefore developed a tool, HHReweighter, utilizing a matrix-element (ME) reweighting method to reduce the total CPU time required to model the full parameter space. The MEs are computed using MadGraph5_amc@nlo (version 2.6.7). Our tool is publicly available in GitLab reweight_repo .
@@ VI.1 Reweighting LO samples
To mitigate this limitation, our strategy is to first decompose the squared ME describing di-Higgs production into several terms corresponding to the three diagrams shown in Fig. 2, as well as the corresponding interference terms. This includes three squared terms corresponding to the $\Box$, $\mathrm{S_{h}}$, and $\mathrm{S_{H}}$ diagrams, plus three cross-terms describing the interferences between these contributions: $-\mathrm{S_{h}}{\text{-}}\Box$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$. The MEs depend on the Yukawa couplings of the $h$ and $H$ to the top quark, denoted $y_{t}^{h}$ and $y_{t}^{H}$, respectively, the Yukawa couplings to the bottom quark, denoted $y_{b}^{h}$ and $y_{b}^{H}$, respectively, and the $\lambda_{hhh}$ and $\lambda_{Hhh}$ trilinear couplings. For convenience, we compute each of the MEs setting $y_{t}^{h}=y_{t}^{H}=y_{t}^{\mathrm{SM}}$, $y_{b}^{h}=y_{b}^{H}=y_{b}^{\mathrm{SM}}$, and $\lambda_{hhh}=\lambda_{Hhh}=\lambda_{\mathrm{SM}}$, where $y_{t}^{\mathrm{SM}}$ ($y_{b}^{\mathrm{SM}}$) and $\lambda_{\mathrm{SM}}$ are the top (bottom) Yukawa and trilinear couplings of the SM Higgs sector. We denote the squared MEs $\mathcal{M}_{i}^{2}$ (where $i=\Box,\mathrm{S_{h}},\mathrm{S_{H}},\mathrm{S_{h}}{\text{-}}\Box,\mathrm{S_{H}}{\text{-}}\Box,\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$) in the following. For the interference terms, we use the shorthand notation
@@
where $\kappa_{q}^{H}\equiv\frac{y_{t}^{H}}{y_{t}^{\mathrm{SM}}}=\frac{y_{b}^{H}}{y_{b}^{\mathrm{SM}}}$, $\kappa_{q}^{h}\equiv\frac{y_{t}^{h}}{y_{t}^{\mathrm{SM}}}=\frac{y_{b}^{h}}{y_{b}^{\mathrm{SM}}}$, and $\kappa_{\lambda_{Hhh}}\equiv\frac{\lambda_{Hhh}}{\lambda_{\mathrm{SM}}}$ [^9] In our current implementation, all quark Yukawa couplings are modified by universal $\kappa_{q}^{h}$ and $\kappa_{q}^{H}$. A generalization to individual Yukawa coupling modifiers will be addressed in future work, e.g. to accommodate also the coupling structure of the 2HDM.. We note that $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms also depend on the assumed values of $m_{H}$ and $\Gamma_{H}$.
The $\mathcal{M}_{i}^{2}$ values can be used to compute weights for the individual components allowing the di-Higgs distributions to be decomposed into the individual terms. An example of this is shown in Fig. 8, for $m_{H}=600\,{\,\rm GeV}$ and $\Gamma_{H}=12\,{\,\rm GeV}$. All distributions are displayed for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$.
We briefly summarize the procedure we used to obtain this figure:
@@
In Fig. 9, we also illustrate the effect of $\Gamma_{H}$ on one of the interference terms, where again contributions have been obtained as described above. All distributions are shown for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{Hhh}}=1$.
Figure 8: The di-Higgs mass distributions split into the individual contributions corresponding to the three diagrams in Fig. 2 and the interference terms. The non-resonant SM-like contributions from the $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$ terms are shown in blue, black, and red, respectively. The $\mathrm{S_{H}}$ resonant contributions is displayed in green. The interference between the $\mathrm{S_{H}}$ and the $\Box$ ($\mathrm{S_{h}}$) diagrams is shown in magenta (cyan). We here use $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$.
@@
  • The physics model in question predicts the values $\kappa_{q}^{h}$, $\kappa_{q}^{H}$, $\kappa_{\lambda_{hhh}}$, $\kappa_{\lambda_{Hhh}}$ and $\Gamma_{H}$ depending on the model parameters. For example, in the singlet model these are computed as a function of the $\tan\beta$, $\sin\alpha$, and $m_{H}$ parameters as
    $\begin{split}\kappa_{q}^{h}&=\cos\alpha,\\ \kappa_{q}^{H}&=\sin\alpha,\\ \kappa_{\lambda_{hhh}}&=\cos^{3}{\alpha}-\tan\beta\sin^{3}{\alpha},\\ \kappa_{\lambda_{Hhh}}&=\frac{2m_{h}^{2}+m_{H}^{2}}{m_{h}^{2}}\,\frac{\sin\left(2\alpha\right)}{2}(\cos{\alpha}+\tan\beta\sin{\alpha})\end{split}$ (29)
@@
Fig. 10 shows examples of the reweighted distributions (blue) compared to the distributions obtained from a MC sample generated directly for the target model parameters (black). The distributions are shown for $m_{hh}$ and $p_{\mathrm{T}}$ of the leading $h$. However, several other variables such as $p_{T}^{h_{1}}$, $p_{T}^{h_{2}}$, $\Delta\eta(h_{1}h_{2})$, $\Delta\phi(h_{1}h_{2})$, $\Delta R(h_{1}h_{2})$, $p_{T}^{hh}$ and $m_{hh}$ with finer bins were analyzed and found to have the same level of agreement.
In this case the target parameter points are $\sin\alpha=0.17$, $\tan\beta=1.5$, and $m_{H}=600\,{\,\rm GeV}$, which yields a predicted width of $5\,{\,\rm GeV}$. We obtain the $\mathrm{S_{H}}$ component by reweighting an $\mathrm{S_{H}}$ MC sample with $\Gamma_{H}=12\,{\,\rm GeV}$ and $\kappa_{\lambda_{Hhh}}=\kappa_{q}^{H}=1$, while a sample of SM di-Higgs events is used to obtain the other terms, as described above. The red line in Fig. 10 shows the $m_{hh}$ spectrum for the MC events used in the reweighting procedure before the reweighting is applied for comparison. The distributions obtained for the reweighted events match the directly generated MC sample within the statistical uncertainties. The method is additionally validated for several other benchmark points as shown in Appendix B.
Figure 10: The reweighting method is validated by comparing reweighted di-Higgs mass (left) and leading $h$ $p_{\mathrm{T}}$ (right) distributions (blue) to the distributions obtained by directly simulating the di-Higgs events for a set of parameter points (black). The red distribution corresponds to the MC events used for the reweighting, before the weights have been applied. On the left, the blue line overlays exactly the black points including error bars.
@@ VI.2 Including higher-order corrections
The discussions so far have focused on MC events generated at LO. However, it is well known that the LO cross-section for gluon-fusion is significantly different compared to higher-order predictions (see e.g. LHCHiggsCrossSectionWorkingGroup:2016ypw and references therein.). It is therefore necessary to correct the LO cross-section through the application of k-factors to account for this. The NNLO cross-sections for the $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$ components deFlorian:2017qfk ; Buchalla:2018yce ; Grazzini:2018bsd ; Amoroso:2020lgh are used to define k-factors for these terms, which we call $k_{\Box}$, $k_{\mathrm{S_{h}}}$, and $-k_{\mathrm{S_{h}}{\text{-}}\Box}$, respectively. For the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ components, no dedicated NNLO predictions are provided. However, NNLO predictions are available for the production of a single undecayed SM-like $H$ ($gg\rightarrow H$LHCHiggsCrossSectionWorkingGroup:2016ypw , and we use these predictions to define the k-factor for the $\mathrm{S_{H}}$ term:
@@
Tab. 3 lists the LO and NNLO cross-sections and the computed k-factors. All cross-sections are computed for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$. The cross-sections displayed for the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms are computed for $m_{H}=600\,{\,\rm GeV}$ and $\Gamma_{H}=5\,{\,\rm GeV}$.
We produce di-Higgs mass distributions for each of the terms in Eq. (28) generated at LO before and after the application of the k-factors. These distributions are shown by the red lines in Figs. 1116. The left and right plots show the distributions before and after the k-factor applications, respectively. We also compare the LO distributions to NLO predictions for cases where MC production is possible in POWHEG, namely the $\Box$, $\mathrm{S_{h}}$, $-\mathrm{S_{h}}{\text{-}}\Box$, and $\mathrm{S_{H}}$ terms. These predictions are shown by the black lines in the figures.
@@
In the POWHEG description Frixione:2007nu , the largest part of the events is simulated with NLO kinematics, giving rise to an additional hard real emission parton [^12] Born type kinematics are Sudhakov supressed. We thank P. Nason for useful comments concerning this topic.. In order to map this to the LO matrix elements used for reweighting, we apply the following procedure: We “ignore” the additional parton when we perform the reweighting. Technically, this is obtained in the following way. We take the two outgoing $h$ four-momenta from the ME and boost to the di-Higgs rest frame. We then obtain the four-momenta for the incoming gluons by requiring both gluons to have zero transverse momentum, and equal and opposite longitudinal momentum, also requiring four-momentum conservation between incoming and outgoing particles. When estimating the MEs we also average over all possible spin-states of the incoming gluons. An NLO SM di-Higgs sample is reweighted to obtain the NLO-approx $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, while an NLO $\mathrm{S_{H}}$ sample with $\Gamma_{H}=12\,{\,\rm GeV}$ is reweighted to obtain the NLO-approx $\mathrm{S_{H}}$ sample for a different width ($\Gamma_{H}=5\,{\,\rm GeV}$ in this case). The NLO-approx results are shown by the blue lines in the figures. The total rate in the NLO-approx method is obtained using the newly determined weights, derived from the LO matrix elements, for each separate contribution and phase space point [^13] In practise this leads to weighted events. We have guaranteed that we sampled the corresponding contributions with sufficient statistical accuracy.. Furthermore, in order to include corrections to the overall rate to NNLO, we define an equivalent set of k-factors accounting for the scaling of the NLO and NLO-approx cross-sections to the NNLO predictions. The NLO(-approx) cross-sections and k-factors are also displayed in Tab. 3.
@@
As expected, Figs. 1113 show a significant difference between the LO and NLO yields due to the differences in the cross-sections. However, after application of the k-factors the yields agree (by construction) and we observe a good agreement in the shapes of the distributions compared to the NLO, indicating that LO samples are adequate for describing the $m_{hh}$ spectrum. The shapes of the NLO-approx distributions also agree well with the NLO distributions, and in this case the yields/cross-section before application of the k-factors are also noticeable quite similar to the NLO results. Figs. 15 and 16 show the LO and NLO-approx distributions for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, respectively. While no comparisons to the full NLO results are possible in this case, the agreement between the shapes and the yields (after k-factor scaling) of the LO and NLO-approx distributions gives further confidence in the description of the $m_{hh}$ spectrum by the LO MC samples after reweighting.
@@
While the LO MC gives a good description of the $m_{hh}$ spectrum, it is well know that the modelling of variables that are sensitive to the additional radiation are more problematic, as this is approximated by the parton shower. In Fig. 17 we show an example of such a variable, the di-Higgs $p_{\mathrm{T}}$, $p_{\mathrm{T}}^{hh}$. The LO samples clearly tend to predict a softer $p_{\mathrm{T}}$ distribution compared to the NLO. In contrast, the agreement of the NLO-approx method is almost exact.
@@
    As shown in Figs. 15 and  16, the predicted $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ yields (and equivalently the cross-sections) for the LO and NLO-approx methods after the k-factor scaling are in good agreement. The cross-sections predicted for the $-\mathrm{S_{H}}{\text{-}}\Box$ term for LO and NLO-approx are 13.52 fb and 13.35 fb, respectively. While the cross-sections for the $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ term are -4.02 fb and -4.00 fb. This demonstrates that the method is, at least, robust regardless of whether it is applied to LO or NLO samples.
@@
    We can apply the same ansatz to estimate the k-factors for the $-\mathrm{S_{h}}{\text{-}}\Box$ term and compare to the k-factors computed directly from the NNLO cross-sections. The k-factors predicted via the ansatz are 2.77, 1.19, and 1.26 for the LO, NLO, and NLO-approx methods, respectively, which agree with the values presented in Tab. 3 within 1%.

### v2

@@ VI Modelling interference effects using reweighting
There are numerous examples of extended Higgs sector models that include an additional scalar that can impact the di-Higgs cross-section and differential distributions, see e.g. Ref. wg3ext for an overview of models currently investigated by the LHC experiments. The final di-Higgs spectrum depends on the parameters of the model. For example, in our investigated singlet model there are three free parameters that affect the relative contributions of the diagrams shown in Figure 2 to the total cross-section, and there are countless other models that may contain even more free parameters. While in principle it is possible to generate Monte Carlo (MC) events for different combinations of parameters, and simulate the detector responses, the computational resources required would be immense. We have therefore developed a tool, HHReweighter, utilising a matrix-element (ME) reweighting method to reduce the total CPU time required to model the full parameter space. The MEs are computed using MadGraph5_amc@nlo (version 2.6.7). Our tool is publicly available in GitLab reweight_repo .
@@ VI.1 Reweighting LO samples
To mitigate this limitation, our strategy is to first decompose the squared ME describing di-Higgs production into several terms corresponding to the three diagrams shown in Figure 2, as well as the corresponding interference terms. This includes three squared terms corresponding to the $\Box$, $\mathrm{S_{h}}$, and $\mathrm{S_{H}}$ diagrams, plus three cross-terms describing the interferences between these contributions: $-\mathrm{S_{h}}{\text{-}}\Box$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$. The MEs depend on the Yukawa couplings of the $h$ and $H$ to the top quark, denoted $y_{t}^{h}$ and $y_{t}^{H}$, respectively, the Yukawa couplings to the bottom quark, denoted $y_{b}^{h}$ and $y_{b}^{H}$, respectively, and the $\lambda_{hhh}$ and $\lambda_{Hhh}$ trilinear couplings. For convenience, we compute each of the MEs setting $y_{t}^{h}=y_{t}^{H}=y_{t}^{\mathrm{SM}}$, $y_{b}^{h}=y_{b}^{H}=y_{b}^{\mathrm{SM}}$, and $\lambda_{hhh}=\lambda_{Hhh}=\lambda_{\mathrm{SM}}$, where $y_{t}^{\mathrm{SM}}$ ($y_{b}^{\mathrm{SM}}$) and $\lambda_{\mathrm{SM}}$ are the top (bottom) Yukawa and trilinear couplings of the SM Higgs sector. We denote the squared MEs $\mathcal{M}_{i}^{2}$ (where $i=\Box,\mathrm{S_{h}},\mathrm{S_{H}},\mathrm{S_{h}}{\text{-}}\Box,\mathrm{S_{H}}{\text{-}}\Box,\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$) in the following. For the interference terms, we use the shorthand notation
@@
where $\kappa_{q}^{H}\equiv\frac{y_{t}^{H}}{y_{t}^{\mathrm{SM}}}=\frac{y_{b}^{H}}{y_{b}^{\mathrm{SM}}}$, $\kappa_{q}^{h}\equiv\frac{y_{t}^{h}}{y_{t}^{\mathrm{SM}}}=\frac{y_{b}^{h}}{y_{b}^{\mathrm{SM}}}$, and $\kappa_{\lambda_{Hhh}}\equiv\frac{\lambda_{Hhh}}{\lambda_{\mathrm{SM}}}$ [^10] In our current implementation, all quark Yukawa couplings are modified by universal $\kappa_{q}^{h}$ and $\kappa_{q}^{H}$. A generalisation to individual Yukawa coupling modifiers will be addressed in future work, e.g. to accommodate also the coupling structure of the two-Higgs-doublet model.. We note that $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms also depend on the assumed values of $m_{H}$ and $\Gamma_{H}$.
The $\mathcal{M}_{i}^{2}$ values can be used to compute weights for the individual components allowing the di-Higgs distributions to be decomposed into the individual terms. An example of this is shown in Figure 8, for $m_{H}=600\,{\,\rm GeV}$ and $\Gamma_{H}=12\,{\,\rm GeV}$. All distributions are displayed for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$.
We briefly summarise the procedure we used to obtain this figure:
@@
In Figure 9, we also illustrate the effect of $\Gamma_{H}$ on one of the interference terms, where again contributions have been obtained as described above. All distributions are shown for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{Hhh}}=1$.
Figure 8: The di-Higgs mass distributions split into the individual contributions corresponding to the three diagrams in Figure 2 and the interference terms. The non-resonant SM-like contributions from the $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$ terms are shown in blue, black, and red, respectively. The $\mathrm{S_{H}}$ resonant contributions is displayed in green. The interference between the $\mathrm{S_{H}}$ and the $\Box$ ($\mathrm{S_{h}}$) diagrams is shown in magenta (cyan). All distributions are shown for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$.
@@
  • The physics model in question predicts the values $\kappa_{q}^{h}$, $\kappa_{q}^{H}$, $\kappa_{\lambda_{hhh}}$, $\kappa_{\lambda_{Hhh}}$, and $\Gamma_{H}$ depending on the model parameters. For example, in the singlet model these are computed as a function of the $\tan\beta$, $\sin\alpha$, and $m_{H}$ parameters as
    $\begin{split}\kappa_{q}^{h}&=\cos\alpha,\\ \kappa_{q}^{H}&=\sin\alpha,\\ \kappa_{\lambda_{hhh}}&=\cos^{3}{\alpha}-\tan\beta\sin^{3}{\alpha},\\ \kappa_{\lambda_{Hhh}}&=\frac{2m_{h}^{2}+m_{H}^{2}}{m_{h}^{2}}\,\frac{\sin\left(2\alpha\right)}{2}(\cos{\alpha}+\tan\beta\sin{\alpha}),\end{split}$ (29)
@@
Figure 10 shows examples of the reweighted distributions (blue) compared to the distributions obtained from a MC sample generated directly for the target model parameters (black). The distributions are shown for $m_{hh}$ and $p_{\mathrm{T}}$ of the leading $h$. However, several other kinematic and angular variables were analysed and found to have the same level of agreement.
In this case the target parameter points are $\sin\alpha=0.17$, $\tan\beta=1.5$, and $m_{H}=600\,{\,\rm GeV}$, which yields a predicted width of $5\,{\,\rm GeV}$. We obtain the $\mathrm{S_{H}}$ component by reweighting an $\mathrm{S_{H}}$ MC sample with $\Gamma_{H}=12\,{\,\rm GeV}$ and $\kappa_{\lambda_{Hhh}}=\kappa_{q}^{H}=1$, while a sample of SM di-Higgs events is used to obtain the other terms, as described above. The red line in Figure 10 shows the $m_{hh}$ spectrum for the MC events used in the reweighting procedure before the reweighting is applied for comparison. The distributions obtained for the reweighted events match the directly generated MC sample within the statistical uncertainties. The method is additionally validated for several other benchmark points as shown in Appendix B.
Figure 10: The reweighting method is validated by comparing the reweighted di-Higgs mass (left) and leading $h$ $p_{\mathrm{T}}$ (right) distributions (blue) to the distributions obtained by directly simulating the di-Higgs events for a set of parameter points (black). The red distribution corresponds to the MC events used for the reweighting, before the weights have been applied. On the left, the blue line overlays exactly the black points including error bars.
@@ VI.2 Including higher-order corrections
The discussions so far have focused on MC events generated at LO. However, it is well known that the LO cross-section for gluon-fusion is significantly different compared to higher-order predictions (see e.g. Ref. LHCHiggsCrossSectionWorkingGroup:2016ypw and references therein). It is therefore necessary to correct the LO cross-section through the application of K-factors to account for this. The NNLO cross-sections for the $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$ components deFlorian:2017qfk ; Buchalla:2018yce ; Grazzini:2018bsd ; Amoroso:2020lgh are used to define K-factors for these terms, which we call $K_{\Box}$, $K_{\mathrm{S_{h}}}$, and $-K_{\mathrm{S_{h}}{\text{-}}\Box}$, respectively. For the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ components, no dedicated NNLO predictions are provided. However, NNLO predictions are available for the production of a single undecayed SM-like $H$ ($gg\rightarrow H$LHCHiggsCrossSectionWorkingGroup:2016ypw , and we use these predictions to define the K-factor for the $\mathrm{S_{H}}$ term:
@@
Table 3 lists the LO and NNLO cross-sections and the computed K-factors. All cross-sections are computed for $\kappa_{q}^{h}=\kappa_{q}^{H}=\kappa_{\lambda_{hhh}}=\kappa_{\lambda_{Hhh}}=1$. The cross-sections displayed for the $\mathrm{S_{H}}$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms are computed for $m_{H}=600\,{\,\rm GeV}$ and $\Gamma_{H}=5\,{\,\rm GeV}$.
We produce di-Higgs mass distributions for each of the terms in Eq. (28) generated at LO before and after the application of the K-factors. These distributions are shown by the red lines in Figures 1116. The left and right plots show the distributions before and after the K-factor applications, respectively. We also compare the LO distributions to NLO predictions for cases where MC production is possible in POWHEG, namely the $\Box$, $\mathrm{S_{h}}$, $-\mathrm{S_{h}}{\text{-}}\Box$, and $\mathrm{S_{H}}$ terms. These predictions are shown by the black lines in the figures.
@@
In the POWHEG description Frixione:2007nu , the largest part of the events is simulated with NLO kinematics, giving rise to an additional hard real emission parton [^13] Born type kinematics are Sudhakov suppressed. We thank P. Nason and E. Re for useful comments concerning this topic.. In order to map this to the LO matrix elements used for reweighting, we apply the following procedure: We “ignore” the additional parton when we perform the reweighting. Technically, this is obtained in the following way. We take the two outgoing $h$ four-momenta from the ME and boost to the di-Higgs rest frame. We then obtain the four-momenta for the incoming gluons by requiring both gluons to have zero transverse momentum, and equal and opposite longitudinal momentum, also requiring four-momentum conservation between incoming and outgoing particles. When estimating the MEs we also average over all possible spin-states of the incoming gluons. An NLO SM di-Higgs sample is reweighted to obtain the NLO-approx $\Box$, $\mathrm{S_{h}}$, and $-\mathrm{S_{h}}{\text{-}}\Box$, $-\mathrm{S_{H}}{\text{-}}\Box$, and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, while an NLO $\mathrm{S_{H}}$ sample with $\Gamma_{H}=12\,{\,\rm GeV}$ is reweighted to obtain the NLO-approx $\mathrm{S_{H}}$ sample for a different width ($\Gamma_{H}=5\,{\,\rm GeV}$ in this case). The NLO-approx results are shown by the blue lines in the figures. The total rate in the NLO-approx method is obtained using the newly determined weights, derived from the LO matrix elements, for each separate contribution and phase space point [^14] In practise this leads to weighted events. We have ensured that we sampled the corresponding contributions with sufficient statistical accuracy.. Furthermore, in order to include corrections to the overall rate to NNLO, we define an equivalent set of K-factors accounting for the scaling of the NLO and NLO-approx cross-sections to the NNLO predictions. The NLO(-approx) cross-sections and K-factors are also displayed in Table 3.
@@
As expected, Figures 1113 show a significant difference between the LO and NLO yields due to the differences in the cross-sections. However, after application of the K-factors the yields agree (by construction) and we observe a good agreement in the shapes of the distributions compared to the NLO, indicating that LO samples are adequate for describing the $m_{hh}$ spectrum. The shapes of the NLO-approx distributions also agree well with the NLO distributions, and in this case the yields/cross-section before application of the K-factors are also noticeable quite similar to the NLO results. Figures 15 and 16 show the LO and NLO-approx distributions for the $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ terms, respectively. While no comparisons to the full NLO results are possible in this case, the agreement between the shapes and the yields (after K-factor scaling) of the LO and NLO-approx distributions gives further confidence in the description of the $m_{hh}$ spectrum by the LO MC samples after reweighting.
@@
While the LO MC gives a good description of the $m_{hh}$ spectrum, it is well know that the modelling of variables that are sensitive to the additional radiation are more problematic, as this is approximated by the parton shower. In Figure 17 we show an example of such a variable, the di-Higgs $p_{\mathrm{T}}$, $p_{\mathrm{T}}^{hh}$. The LO samples clearly tend to predict a softer $p_{\mathrm{T}}$ distribution compared to the NLO. In contrast, the agreement of the NLO-approx method is almost exact.
@@
    As shown in Figures 15 and  16, the predicted $-\mathrm{S_{H}}{\text{-}}\Box$ and $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ yields (and equivalently the cross-sections) for the LO and NLO-approx methods after the K-factor scaling are in good agreement. The cross-sections predicted for the $-\mathrm{S_{H}}{\text{-}}\Box$ term for LO and NLO-approx are 13.52 fb and 13.35 fb, respectively. While the cross-sections for the $\mathrm{S_{H}}{\text{-}}\mathrm{S_{h}}$ term are -4.02 fb and -4.00 fb. This demonstrates that the method is, at least, robust regardless of whether it is applied to LO or NLO samples.
@@
    We can apply the same ansatz to estimate the K-factors for the $-\mathrm{S_{h}}{\text{-}}\Box$ term and compare to the K-factors computed directly from the NNLO cross-sections. The K-factors predicted via the ansatz are 2.77, 1.19, and 1.26 for the LO, NLO, and NLO-approx methods, respectively, which agree with the values presented in Table 3 within 1%.